### v2

# ELASTIC: Efficient Once For All Iterative Search for Object Detection on Microcontrollers

Abstract
Deploying high-performance object detectors on TinyML platforms poses significant challenges due to tight hardware constraints and the modular complexity of modern detection pipelines. Neural Architecture Search (NAS) offers a path toward automation, but existing methods either restrict optimization to individual modules—sacrificing cross-module synergy—or require global searches that are computationally intractable. We propose ELASTIC (Efficient Once for AlL IterAtive Search for ObjecT DetectIon on MiCrocontrollers), a unified, hardware-aware NAS framework that alternates optimization across modules (e.g., backbone, neck, and head) in a cyclic fashion. ELASTIC introduces a novel Population Passthrough mechanism in evolutionary search that retains high-quality candidates between search stages, yielding faster convergence, up to an 8% final mAP gain, and eliminates search instability observed without population passthrough. In a controlled comparison, empirical results show ELASTIC achieves +4.75% higher mAP and 2× faster convergence than progressive NAS strategies on SVHN, and delivers a +9.09% mAP improvement on PascalVOC given the same search budget. ELASTIC achieves 72.3% mAP on PascalVOC, outperforming MCUNET by 20.9% and TinyissimoYOLO by 16.3%. When deployed on MAX78000/MAX78002 microcontrollers, ELASTIC-derived models outperform Analog Devices’ TinySSD baselines, reducing energy by up to 71.6 %, lowering latency by up to 2.4, and improving mAP by up to 6.99 percentage points across multiple datasets.
I Introduction
Object detection is a fundamental task in computer vision with real-world applications in mobile robotics, autonomous vehicles, smart surveillance, and embedded systems [1, 2]. While modern detectors achieve remarkable accuracy and speed on powerful hardware, their deployment on TinyML platforms—microcontrollers and edge devices with sub-MB memory, ultra-low power, and limited compute budgets—remains challenging [3, 4, 5, 6, 7].
However, unlike image classification, object detection involves multi-stage pipelines—typically composed of a backbone [8], neck [9], and head [1, 2]—each contributing uniquely to accuracy, latency, and memory footprint [10]. These interdependent components enlarge the design space exponentially, complicating deployment on microcontrollers, where model size, SRAM usage, and inference cycles must be tightly controlled [11, 5, 12].
Neural Architecture Search (NAS) offers a compelling approach to automate architecture design under tight resource constraints [13, 14]. Constraint-aware NAS has advanced edge classification [15, 3, 16], but object detection adds new layers of complexity. Prior detection NAS methods focus on individual modules (e.g., backbones [17], necks [9], or heads [18]), often ignoring cross-module dependencies. Full-pipeline NAS is possible [19], but scales poorly due to the massive combinatorial space—e.g., over configurations in realistic search settings [18, 20].
To address these limitations, we introduce ELASTIC (Efficient Once for AlL IterAtive Search for ObjecT DetectIon on MiCrocontrollers), a novel once-for-all NAS method that performs constrained iterative search. ELASTIC alternates between optimizing one module while fixing others in a cyclic fashion, leveraging previous iteration feedback to refine architecture choices. This strategy preserves cross-module interdependencies while dramatically reducing the effective search space. Further, our approach enforces explicit resource constraints—such as model size—throughout the search process, ensuring deployability on MCUs.
Our key contributions are summarized as follows:
Unified Iterative NAS: We propose a novel NAS framework that cyclically optimizes different modules (i.e. backbone, neck, head), capturing inter-module interactions and expanding the search space. Experimental results demonstrate that our approach achieves architectures with up to 4.75% higher mAP than progressive search strategies in early stages while being twice as fast.
Population Passthrough Mechanism: We develop a memory-based population retention scheme that mitigates performance drops during module alternation. On PascalVOC, this mechanism improves final mAP from 22.1% without passthrough to 30.83% and stabilizes search convergence across iterations.
Enhanced Search Space: ELASTIC significantly enhances the search space over time. On SVHN, the mean mAP of randomly sampled architectures improves from 44.61% (joint) to 70.68% after 5 iterations, while variance reduces by 89.7%. On PascalVOC, the mean mAP improves from 4.99% to 28.32%, with a 95.3% variance reduction.
Once-for-All Deployment on Resource-Constrained Devices: We validate ELASTIC in designing versatile models tailored to various resource constraints, demonstrating their inference efficiency on the MAX78000 platform. Our ELASTIC model significantly outperforms the baseline [21] on the MAX78000 platform, achieving a 45.4% reduction in energy consumption, a 29.3% decrease in latency, and a 4.5% improvement in mAP, demonstrating enhanced computational efficiency and accuracy.
II Related Work
NAS for Object Detection. Object detection architectures are modular, comprising a backbone, neck, and head [2, 1, 10]. NAS methods have targeted these components individually or jointly: DetNAS [17] searches backbone architectures via a one-shot supernet; NAS-FPN [9] focuses on FPN design; NAS-FCOS [18] extends search to the FPN and head using reinforcement learning; MobileDets [22] introduces latency-aware backbone optimization for mobile accelerators; and Hit-Detector [19] jointly searches all modules but incurs high computational cost. While effective on larger systems, these methods are not designed for deployment on resource-constrained microcontrollers.
Progressive vs. Global Search. Global NAS jointly optimizes all modules but is intractable at TinyML scales due to the combinatorial size of the joint search space [19, 18]. Progressive NAS reduces complexity by optimizing modules sequentially—typically starting with the backbone or neck—while keeping previously optimized components fixed [18, 23]. However, early architectural choices (e.g., output resolution, feature dimensionality, or stage depths) impose rigid interface constraints that limit the feasible design space for subsequent modules, often leading to suboptimal detection accuracy or violating tight memory and compute budgets in later stages [9]. Our approach addresses this by iteratively alternating module optimization, enabling cross-stage adaptation while preserving search efficiency.
OFA and Iterative NAS. Once-for-all (OFA) methods train a supernet capable of representing many sub-networks, enabling fast specialization to various hardware or latency constraints without retraining [15]. DetOFA [20] adapts this strategy to object detection using constraint-aware path pruning. These methods provide a foundation for efficient, hardware-aware NAS; our work builds on this by incorporating modular search dynamics specific to detection pipelines.
III Methodology
We propose ELASTIC, a novel constrained NAS strategy that iteratively optimizes individual modules while maintaining inter-module context. This approach balances search efficiency with architectural synergy, enabling the discovery of compact, high-performing object detectors for real-time deployment on microcontrollers. To further improve convergence, we introduce a Population Passthrough mechanism that preserves top-performing architecture candidates across iterations during evolutionary search.
III-A Problem Formulation
Let an object detection subnet be represented as extracted from a supernet with architecture and weights . The search space contains all valid architectures such that . The objective of NAS is to find the optimal architecture that minimizes the validation loss when evaluated with shared weights :
Here, is the hardware-constrained model size. Inspired by Once-for-All NAS [15], a weight-shared supernet reduces training cost. Since grows combinatorially for joint optimization (e.g., ), optimization alternates between backbone and head . To extend this formulation to neck modules, the neck design choices can be integrated into the head search space, enabling joint optimization of both neck and head components within a single search stage.
At iteration , we first optimize the backbone while keeping fixed:
Then, we fix and update the head:
We enforce the constraint , where is the global memory budget. This formulation ensures that each module’s optimization phase respects deployment constraints.
III-B Iterative Evolutionary Architecture Search
At the core of our ELASTIC framework is an Iterative Evolutionary Architecture Search, which alternates optimization between modular components—specifically the backbone and the head—over successive cycles. Unlike conventional evolutionary NAS that operates in a single global pass, our method applies modular evolutionary search iteratively, enabling progressive refinement and co-adaptation across modules. This design explicitly captures cross-module interactions while maintaining search tractability under TinyML constraints.
The evolutionary procedure within each iteration follows these steps:
Population Initialization: The candidate population is composed of two parts: (1) a set of high-performing architectures retained from the module’s previous iteration via Population Passthrough, and (2) newly sampled architectures that maintain diversity in the search.
Genetic Variation. We generate new architectural candidates using crossover and mutation, subject to module-specific resource constraints (e.g., for the backbone and for the head). All offspring are filtered to ensure constraints are satisfied.
Fitness Evaluation: Each candidate is instantiated within the weight-shared supernet and evaluated using mean average precision (mAP) on a validation set. Since only one module is varied at a time, the evaluation reflects the candidate’s performance in the context of a fixed complementary module.
Selection: The highest-performing candidates are retained to seed the next generation. This ensures that architectural refinements are preserved and that future iterations start from strong initial configurations.
After optimizing the current module (e.g., backbone), the algorithm switches focus to the other module (e.g., head) and repeats the process. This iterative back-and-forth search continues for a predefined number of cycles or until convergence criteria are met. This structured alternation between modules allows ELASTIC to gradually co-adapt modular components under deployment constraints.
III-C Population Passthrough
A central challenge in iterative modular NAS is maintaining optimization continuity across alternating search phases. When a search process shifts focus from one module (e.g., backbone) to another (e.g., head), conventional evolutionary algorithms often discard the entire population from the previous cycle. This reinitialization disrupts search momentum, leading to slower convergence and redundant exploration.
| Dataset | Method | mAP | mAP | Cost | MACs | Params | Latency |
| SVHN | Backbone-Only [17] | 75.53% | +0.18% | 10.8 hrs | 137.6 M | 0.73 M | 2.62 ms |
| Head-Only [18] | 75.35% | +0.00% | 4.7 hrs | 475.3 M | 1.01 M | 2.69 ms | |
| Progressive [18] | 79.62% | +4.27% | 30.8 hrs | 78.4 M | 0.52 M | 2.36 ms | |
| ELASTIC (OURS) | 80.09% | +4.74% | 12.5 hrs | 105.8 M | 0.61 M | 2.34 ms | |
| PascalVOC | Backbone-Only [17] | 22.02% | +2.80% | 9.6 hrs | 436.1 M | 0.58 M | 2.11 ms |
| Head-Only [18] | 19.22% | +0.00% | 6.0 hrs | 240.0 M | 0.36 M | 2.48 ms | |
| Progressive [18] | 21.74% | +2.52% | 14.8 hrs | 540.0 M | 0.41 M | 2.66 ms | |
| ELASTIC (OURS) | 30.83% | +11.61% | 14.65 hrs | 642.8 M | 0.57 M | 2.00 ms |
| Method | MACs | MACs | Params | VOC mAP | mAP |
| TY: 20-3-88 [24] | 32M | 90.7% | 0.58M | 53% | +30% |
| TY: 20-7-88 [24] | 44M | 87.2% | 0.58M | 47% | +24% |
| TY: 20-3-112 [24] | 54M | 84.3% | 0.89M | 56% | +33% |
| TY: 20-7-112 [24] | 70M | 79.6% | 0.91M | 53% | +30% |
| TY: 20-3-224 [24] | 218M | 36.4% | 3.34M | 23% | +0% |
| MCUNet [3] | 168M | 51.0% | 1.2M | 51.4% | +28.4% |
| MCUNetV2-M4 [25] | 172M | 49.9% | 0.47M | 64.6% | +41.6% |
| MCUNetV2-H7 [25] | 343M | 0% | 0.67M | 68.3% | +45.3% |
| ELASTIC (OURS) | 86M | 74.9% | 1.36M | 72.3% | +49.3% |
To mitigate this, we introduce Population Passthrough, a memory-based mechanism that explicitly retains high-quality search knowledge across module alternations. Each module maintains a passthrough memory buffer , which stores a ranked set of elite candidates based on performance from its most recent optimization cycle. When control returns to module in a later iteration, the initial population for evolutionary search is composed as follows:
Elite Inheritance: A partition of the population is sampled directly from , preserving high-performing architectural patterns discovered in earlier cycles.
Exploration: The remaining candidates are randomly sampled from the current search space, ensuring sufficient diversity to explore new regions and avoid premature convergence.
Notably, Population Passthrough differs from traditional aging [14], which typically operate within a single module or generation. Instead, it functions across alternating optimization cycles in a modular setting, making it uniquely suited for constrained multi-part architectures such as those required by TinyML.
III-D Resource Allocation and Search Budget
To ensure deployability on resource-constrained hardware, we explicitly partition the overall memory budget between architectural modules. Let and denote the memory allocations for the backbone and head, respectively, such that . This modular budget decomposition is critical for embedded systems, where hardware constraints—such as on-chip SRAM, flash storage, and runtime latency—are rigid and non-negotiable [6]. By enforcing per-module constraints during search, ELASTIC ensures that all discovered subnetworks are compatible with the deployment requirements of target TinyML platforms.
In addition, ELASTIC is designed to support search interruption, allowing the iterative process to be paused at any stage and a valid, high-quality subnetwork to be extracted. This property is particularly beneficial in real-world TinyML development workflows, where compute resources are limited and iterative refinement may need to occur incrementally [4]. The ability to recover performant models mid-search adds flexibility and practicality to constrained NAS pipelines.
IV Experiments
| Dataset | Search Space | Mean mAP | mAP | Std. Dev. | Variance |
| SVHN | Joint Search [19] | 44.61% | 0% | 1.39 | |
| Head Search [18] | 67.38% | +22.77% | 0.37 | 0.136 | |
| ELASTIC(3) | 70.42% | +25.81% | 0.47 | 0.221 | |
| ELASTIC(5) | 70.68% | +26.07% | 0.45 | 0.199 | |
| PascalVOC | Joint Search [19] | 4.99% | 0% | 0.04 | |
| Head Search [18] | 27.89% | +22.90% | 0.01 | ||
| ELASTIC(3) | 28.28% | +23.29% | 0.01 | ||
| ELASTIC(5) | 28.32% | +23.33% | 0.01 |
IV-A Experimental Setup
We construct four OFA-based supernets tailored for diverse TinyML-relevant datasets: (i) SVHN, (ii) VGGFace2, (iii) a five-class subset of PascalVOC, and (iv) the full PascalVOC dataset. These datasets are selected to cover a range of TinyML application domains, from lightweight digit recognition to face identification and general-purpose object detection, while avoiding the prohibitive memory and compute demands of large-scale datasets such as COCO.
The SVHN, VGGFace2, and PascalVOC subset supernets adopt a one-stage SSD-style architecture, while the full PascalVOC supernet adopts a MobileNet–YOLO style design that combines a lightweight MobileNet backbone [8] with YOLO-style detection heads [1] to accommodate the increased dataset complexity. For the SSD-style supernets, the search space varies (i) layer widths, (ii) kernel sizes (, ), and (iii) depth, i.e., the number of layers allocated at each feature map. For the MobileNet–YOLO supernet, the search space varies (i) expansion ratios in inverted bottlenecks, (ii) kernel sizes (, ), and (iii) depth, i.e., the number of inverted bottleneck blocks per feature map.
All supernets are trained with a batch size of 64, learning rate of 0.001, momentum of 0.9, and weight decay of . Training follows the progressive shrinking strategy [15], with 200 epochs per level transition. Experiments are conducted on a single NVIDIA RTX 4090 GPU.
Evolutionary search is performed with a population size of 100, mutation probability of 0.2, mutation ratio of 0.5, and a parent selection ratio of 0.25. Each search is budgeted for 55 generations on SVHN and VGGFace2, and 30 generations on the PascalVOC subset and full PascalVOC datasets, respectively. Convergence is defined as the point at which the best candidate architecture’s mAP is within 1% of the final value achieved at the end of the search.
IV-B Comparative Search Evaluation
We evaluate the effectiveness of ELASTIC by comparing it against three baseline NAS strategies: backbone-only search [17], head-only search [18], and progressive NAS [18]. All methods are evaluated under a fixed budget using a consistent once-for-all supernet design.
Results on SVHN and PascalVOC subset.
Table I summarizes our search results. On SVHN, ELASTIC achieves a final mAP of 80.09%, outperforming progressive NAS (79.62%) while using only 12.5 GPU hours—a 2 speedup compared to the 30.8 hours required by progressive search. On PascalVOC, the improvement is more substantial: ELASTIC reaches 30.83% mAP, outperforming progressive NAS (21.74%) by +9.09 percentage points with similar compute (14.65 vs. 14.78 GPU hours).
Effect of Decoupling.
To isolate the impact of coordinated multi-module optimization, we compare ELASTIC against decoupled strategies that search only the backbone or the head. On SVHN, these methods achieve significantly lower performance: backbone-only search yields 75.53% mAP and head-only search reaches just 75.35%, falling short of ELASTIC’s 80.09% by 4.75 percentage points.
Convergence Behavior.
Figure 1 shows that ELASTIC converges up to 2 faster than progressive NAS. On SVHN, ELASTIC converges in just 15 iterations, compared to over 30 iterations for progressive NAS. In contrast, backbone-only and head-only searches stagnate early, plateauing at 75.53% and 75.35% mAP, respectively. A similar trend holds on PascalVOC, where ELASTIC rapidly reaches 30.83% mAP, outperforming progressive NAS (21.74%) and single-module baselines (22.02% and 19.22%). This disparity is exacerbated by progressive NAS’s rigid search pipeline, which fixes early module choices and prevents adaptation to downstream module configurations.
Comparison with MCUNET and TinyissimoYOLO.
To assess ELASTIC’s suitability for microcontrollers, we extend our experiments using the MCUNET [3] search space with the neck module included for fair comparison. Benchmarked on the full PascalVOC () dataset with similar parameter budgets (Table II), our ELASTIC-derived detector achieves 72.3% mAP, surpassing MCUNET by 20.9%, TinyissimoYOLO’s [5] best model by 16.3%, and MCUNetV2-M4/H7 [25] by 7.7% and 4.0%, respectively, while using 50–75% fewer MACs. Unlike MCUNetV2 [25], which requires external patch-based inference, ELASTIC supports native end-to-end inference, making it more general and easier to deploy across diverse microcontrollers.
| Model | Dataset | Device | Params | Energy (J) | Latency (ms) | Power (mW) | mAP (%) |
| ai87-fpndetector [21] | PascalVOC | MAX78002 | 2.18M | 62001 | 122.6 | 445.76 | 50.66 |
| ELASTIC (OURS) | PascalVOC | MAX78002 | 1.32M | 17581 | 51.1 | 285.02 | 57.65 |
| ai85net-tinierssd-face [21] | VGGFace2 | MAX78000 | 0.28M | 1712 | 43.4 | 29.90 | 84.72 |
| ELASTIC (OURS) | VGGFace2 | MAX78000 | 0.22M | 1368 | 45.6 | 20.90 | 87.10 |
| ai85net-tinierssd [21] | SVHN | MAX78000 | 0.19M | 573 | 14.0 | 29.20 | 83.60 |
| ELASTIC (OURS) | SVHN | MAX78000 | 0.22M | 341 | 13.0 | 16.70 | 88.10 |
IV-C Impact of Population Passthrough on Search Stability and Efficiency
Iterative NAS often suffers from performance drops when switching between modules due to population reinitialization. Without memory, each alternation incurs a “catch-up” phase, degrading search stability and slowing convergence. To address this, we introduce Population Passthrough, a memory mechanism that retains top-performing architectures across module alternations.
Figure 3 shows the effect of varying passthrough ratios on the PascalVOC dataset. Without passthrough, mAP repeatedly drops after each module switch and saturates below 22%. In contrast, ratios of 50–80% yield smoother trajectories and reach over 30% mAP within half the iterations. The best results are obtained with a 60% passthrough ratio, which strikes an effective balance between inheritance of strong candidates and exploration of new architectures. Higher ratios risk overfitting to stale subnets, while lower ratios lead to instability.
IV-D Iterative Search Space Refinement
To evaluate how ELASTIC improves the underlying search space over time, we conduct a series of random sampling experiments across both SVHN and PascalVOC. Specifically, we compare the distribution of model accuracies under three search conditions: (1) the original joint search space, (2) the progressive head-only search space, and (3) the refined head space produced by ELASTIC after iterative search.
Figure 4 shows the accuracy scatter plots and means of 100 randomly sampled architectures from each space. On SVHN, the joint space produces models with a mean accuracy of 44.61%, while progressive head-only search improves this to 67.38%. ELASTIC further enhances the head space, achieving a mean accuracy of 72.25% — a 4.87% increase over progressive search. A similar trend is observed on PascalVOC, where the mean accuracy improves from 4.99% (joint) to 27.89% (progressive), and then to 28.32% with ELASTIC.
To further examine the progression of search space quality, we track mean and variance over time using random samples taken at key iterations of ELASTIC. As shown in Figure 5 and Table III, the mean accuracy on SVHN improves from 67.38% at iteration 0 to 70.68% at iteration 5, while variance decreases, indicating a more stable and optimized search space. Similar patterns are observed on PascalVOC, where the mean increases from 27.89% to 28.32%, with variance decreasing from to .
IV-E Deployment on MAX78000/MAX78002 Platform
The MAX78000/MAX78002 platform integrates a dedicated CNN accelerator but imposes strict architectural and memory constraints. For example, the MAX78000 includes a 432 KB weight memory limit, 32 KB of data memory, support restricted to and convolution kernels, a maximum of 32 layers, and up to 1024 channels per layer while the MAX78002 includes a 2.3 MB weight memory limit, 80 KB of data memory, and a maximum of 128 layers.
Despite these limitations, our ELASTIC-derived models demonstrate substantial improvements over Analog Devices’ baselines [21]. On PascalVOC, ELASTIC reduces energy by 71.6%, lowers latency by 2.40, and improves mAP by +6.99, while remaining within the on-chip memory budget. On VGGFace2 and SVHN, ELASTIC achieves consistent efficiency gains, reducing energy by 20.1% and 40.5%, power by 30.1% and 42.8%, and latency by 5.1%and 7.1%, respectively, with mAP gains of +2.38% and +4.50%.
Figure 6 compares ELASTIC-derived architectures with manually scaled TinierSSD [21] baselines in terms of accuracy and FLOPs. ELASTIC consistently achieves higher accuracy while significantly reducing FLOPs—up to 66%— at similarly sized models, highlighting the benefits of hardware-aware NAS.
V Conclusion
We introduced ELASTIC, a unified framework for constrained, once-for-all NAS tailored to object detection on resource-limited platforms. By alternating backbone and head optimization, ELASTIC enables cross-module co-adaptation while significantly reducing search cost. A key component of our approach, Population Passthrough, stabilizes module transitions and accelerates convergence, addressing a major limitation of existing progressive and decoupled NAS methods.
Through extensive experiments on multiple datasets and deployment on real microcontroller platforms, we demonstrated that ELASTIC discovers architectures that are both more accurate and more efficient than manually designed and standard NAS baselines. These results highlight the effectiveness of iterative, hardware-aware NAS in advancing the accuracy–efficiency frontier for resource-constrained edge devices.