### v3

# ELASTIC: Efficient Once For All Iterative Search for Object Detection on Microcontrollers

Abstract
Deploying high-performance object detectors on TinyML platforms poses significant challenges due to tight hardware constraints and the modular complexity of modern detection pipelines. Neural Architecture Search (NAS) offers a path toward automation, but existing methods either restrict optimization to individual modules—sacrificing cross-module synergy—or require global searches that are computationally intractable. We propose ELASTIC (Efficient Once for AlL IterAtive Search for ObjecT DetectIon on MiCrocontrollers), a unified, hardware-aware NAS framework that alternates optimization across modules (e.g., backbone, neck, and head) in a cyclic fashion. ELASTIC introduces a novel Population Passthrough mechanism in evolutionary search that retains high-quality candidates between search stages, yielding faster convergence, up to an 8% final mAP gain, and eliminates search instability observed without population passthrough. In a controlled comparison, empirical results show ELASTIC achieves +4.75% higher mAP and 2× faster convergence than progressive NAS strategies on SVHN, and delivers a +9.09% mAP improvement on PascalVOC given the same search budget. ELASTIC achieves 72.3% mAP on PascalVOC, outperforming MCUNET by 20.9% and TinyissimoYOLO by 16.3%. When deployed on MAX78000/MAX78002 microcontrollers, ELASTIC-derived models outperform Analog Devices’ TinySSD baselines, reducing energy by up to 71.6 %, lowering latency by up to 2.4, and improving mAP by up to 6.99 percentage points across multiple datasets. The experimental videos and codes are available on the project website111Project Website: https://nail-uh.github.io/elastic.github.io/.
I Introduction
Object detection is a fundamental task in computer vision with real-world applications in mobile robotics, autonomous vehicles, smart surveillance, and embedded systems. While modern detectors achieve remarkable accuracy and speed on powerful hardware, their deployment on TinyML platforms—microcontrollers and edge devices with sub-MB memory, ultra-low power, and limited compute budgets—remains challenging [13, 11, 15, 17, 25].
Unlike image classification, object detection relies on a multi-stage pipeline—backbone, neck [7], and head—whose design choices are strongly interdependent: modifying one module changes the feature distributions and optimal configurations of the others. This cross-module coupling makes the detector design space highly combinatorial, which is particularly challenging on microcontrollers where flash/SRAM budgets and inference cycles must be tightly controlled [21, 15, 1]. Recent hardware-aware NAS efforts for pure MCU targets have largely focused on image classification [5, 6], where the optimization is over single-stream networks with a single prediction head and classification objectives; in contrast, MCU detection requires jointly reasoning over the coupled backbone–neck–head pipeline and mAP-driven evaluation.
Recent TinyML-oriented detection research improves efficiency via specialized architectures and deployment-aware design. Hardware-conscious model and compilation/runtime co-design enable practical MCU execution [13, 1, 21], while lightweight detectors and variants (including transformer-inspired designs) explore accuracy–efficiency trade-offs under tight energy/latency constraints [15, 17, 25, 11]. Despite these advances, selecting an optimal detector for a specific MCU remains non-trivial because accuracy, flash/SRAM footprint, and inference cycles are tightly coupled, and manual design or ad-hoc scaling rarely navigates these constraints systematically at the full detector level.
Neural Architecture Search (NAS) offers a compelling approach to automate architecture design under tight resource constraints. Constraint-aware NAS has advanced edge classification [2, 13, 14], but object detection adds new layers of complexity. Prior detection NAS methods focus on individual modules (e.g., backbones [4], necks [7], or heads [22]), often ignoring cross-module dependencies. Full-pipeline NAS is possible [8], but scales poorly due to the massive combinatorial space—e.g., over configurations in realistic search settings [22, 19]. These observations expose a key gap: MCU-targeted detection NAS must preserve cross-module co-adaptation (backbone–neck–head interactions) without incurring the prohibitive cost of full end-to-end search, while enforcing hard resource constraints so discovered models are directly deployable.
To address these limitations, we introduce ELASTIC (Efficient Once for AlL IterAtive Search for ObjecT DetectIon on MiCrocontrollers), a novel once-for-all NAS method that performs constrained iterative search. ELASTIC alternates between optimizing one module while fixing others in a cyclic fashion, leveraging previous iteration feedback to refine architecture choices. This strategy preserves cross-module interdependencies while dramatically reducing the effective search space. Further, our approach enforces explicit resource constraints—such as model size—throughout the search process, ensuring deployability on MCUs.
Our key contributions are summarized as follows:
Unified Iterative NAS: We propose a cyclic module-wise NAS framework that alternates optimization across the backbone, neck, and head to preserve cross-module co-adaptation while reducing search cost by restricting the active search dimension at each iteration. In our early-stage PascalVOC search, ELASTIC achieves up to +4.75% mAP over the progressive baseline while on SVHN, ELASTIC reduces search cost until convergence from 30.8 to 12.5 GPU hours.
Population Passthrough Mechanism: We introduce Population Passthrough, an elite-carryover strategy that stabilizes the iterative module alternation by retaining top-performing candidates when switching the active search module, rather than reinitializing the evolutionary population. This directly mitigates the abrupt performance drops observed under naive module switching and enables consistent convergence across iterations. On PascalVOC, enabling passthrough improves the final mAP from 22.1% (iterative search without passthrough) to 30.83% (with passthrough), demonstrating that the mechanism is a necessary component for making cyclic multi-module optimization effective in practice.
Enhanced Search Space: ELASTIC enhances the search space over time. On SVHN, the mean mAP of randomly sampled architectures improves from 44.61% (joint) to 70.68% after 5 iterations, while variance reduces by 89.7%. On PascalVOC, the mean mAP improves from 4.99% to 28.32%, with a 95.3% variance reduction.
Once-for-All Deployment on Resource-Constrained Devices: We validate ELASTIC in designing versatile models tailored to various resource constraints, demonstrating their inference efficiency on the MAX78000/02 and STM32F746 platform. Our ELASTIC model significantly outperforms the baseline [9] on the MAX78000 platform, achieving a 45.4% reduction in energy consumption, a 29.3% decrease in latency, and a 4.5% improvement in mAP, demonstrating enhanced computational efficiency and accuracy.
II Related Work
NAS for Object Detection. Object detection architectures are modular, comprising a backbone, neck, and head. NAS methods have targeted these components individually or jointly: DetNAS [4] searches backbone architectures via a one-shot supernet; NAS-FPN [7] focuses on FPN design; NAS-FCOS [22] extends search to the FPN and head using reinforcement learning; MobileDets [23] introduces latency-aware backbone optimization for mobile accelerators; and Hit-Detector [8] jointly searches all modules but incurs high computational cost. While effective on larger systems, these methods are not designed for deployment on resource-constrained microcontrollers.
Progressive vs. Global Search. Global NAS jointly optimizes all modules but is intractable at TinyML scales due to the combinatorial size of the joint search space [8, 22]. Progressive NAS reduces complexity by optimizing modules sequentially—typically starting with the backbone or neck—while keeping previously optimized components fixed [22, 18]. However, early architectural choices (e.g., output resolution, feature dimensionality, or stage depths) impose rigid interface constraints that limit the feasible design space for subsequent modules, often leading to suboptimal detection accuracy or violating tight memory and compute budgets in later stages [7]. Our approach addresses this by iteratively alternating module optimization, enabling cross-stage adaptation while preserving search efficiency.
OFA and Iterative NAS. Once-for-all (OFA) methods train a supernet capable of representing many sub-networks, enabling fast specialization to various hardware or latency constraints without retraining [2]. DetOFA [19] adapts this strategy to object detection using constraint-aware path pruning. These methods provide a foundation for efficient, hardware-aware NAS; our work builds on this by incorporating modular search dynamics specific to detection pipelines.
III Methodology
This section introduces ELASTIC, a constrained iterative NAS method for TinyML object detection that explicitly treats the detector as a multi-module pipeline (backbone–neck–head). The key distinction from prior NAS strategies is that ELASTIC performs cyclic module-wise optimization: at each iteration, we optimize one module while keeping the others fixed, and we revisit modules multiple times to enable cross-module co-adaptation without incurring the prohibitive cost of full end-to-end search. However, naively switching the optimized module can destabilize the search and cause abrupt drops in validation performance. To address this, we propose Population Passthrough, which retains top-performing architecture candidates at the end of each module stage and injects them into the next stage, stabilizing convergence across iterations under explicit resource constraints.
III-A Problem Formulation
Let an object detection subnet be represented as extracted from a supernet with architecture and weights . The search space contains all valid architectures such that . The objective of NAS is to find the optimal architecture that minimizes the validation loss when evaluated with shared weights :
Here, is the hardware-constrained model size. Inspired by Once-for-All NAS [2], a weight-shared supernet reduces training cost. Since grows combinatorially for joint optimization (e.g., ), optimization alternates between backbone and head . To extend this formulation to neck modules, the neck design choices can be integrated into the head search space, enabling joint optimization of both neck and head components within a single search stage.
At iteration , we first optimize the backbone while keeping fixed:
Then, we fix and update the head:
We enforce the constraint , where is the global memory budget. This formulation ensures that each module’s optimization phase respects deployment constraints.
III-B Iterative Evolutionary Architecture Search
At the core of ELASTIC is an Iterative Evolutionary Architecture Search that alternates optimization between the backbone and the head over successive cycles. As illustrated in Fig. 1, the framework does not search the full detector space in a single pass. Instead, it fixes one module while optimizing the other, then switches roles and repeats this process iteratively. This design preserves search tractability for TinyML deployment while allowing the two modules to progressively co-adapt across iterations.
Within each stage shown in Fig. 1, the evolutionary search follows a standard population-based procedure, but is applied only to the currently active module. The population is initialized using two sources: high-performing candidates carried over from the module’s previous visit through Population Passthrough, and newly sampled candidates that maintain exploration. New architectures are then generated through crossover and mutation under the corresponding resource constraints (e.g., for the backbone and for the head). Each candidate is evaluated inside the weight-shared supernet using validation mAP, where the complementary module remains fixed. The best-performing candidates are retained to guide the next generation and to provide a strong starting point when the search returns to that module in a later cycle.
As depicted in Fig. 1, once the search for the current module is completed, ELASTIC switches to the other module and repeats the same procedure. By alternating in this back-and-forth manner, the framework incrementally refines both modules rather than optimizing them only once. This repeated revisiting is important because improving one module changes the context in which the other operates. Therefore, the iterative structure of ELASTIC enables the backbone and head to adapt to each other over time while remaining within deployment constraints.
III-C Population Passthrough
| Dataset | Method | mAP | mAP | Cost | MACs | Params | Latency |
| SVHN | Backbone-Only [4, 3, 20] | 75.53% | +0.18% | 10.8 hrs | 137.6 M | 0.73 M | 2.62 ms |
| Neck/Head-Only [7, 10, 24] | 75.35% | +0.00% | 4.7 hrs | 475.3 M | 1.01 M | 2.69 ms | |
| Progressive [22] | 79.62% | +4.27% | 30.8 hrs | 78.4 M | 0.52 M | 2.36 ms | |
| ELASTIC (OURS) | 80.09% | +4.74% | 12.5 hrs | 105.8 M | 0.61 M | 2.34 ms | |
| PascalVOC | Backbone-Only [4] | 22.02% | +2.80% | 9.6 hrs | 436.1 M | 0.58 M | 2.11 ms |
| Neck/Head-Only [7, 10, 24] | 19.22% | +0.00% | 6.0 hrs | 240.0 M | 0.36 M | 2.48 ms | |
| Progressive [22] | 21.74% | +2.52% | 14.8 hrs | 540.0 M | 0.41 M | 2.66 ms | |
| ELASTIC (OURS) | 30.83% | +11.61% | 14.65 hrs | 642.8 M | 0.57 M | 2.00 ms |
| Method | MACs | MACs | Params | VOC mAP | mAP |
| TY: 20-3-88 [16] | 32M | 90.7% | 0.58M | 53% | +30% |
| TY: 20-7-88 [16] | 44M | 87.2% | 0.58M | 47% | +24% |
| TY: 20-3-112 [16] | 54M | 84.3% | 0.89M | 56% | +33% |
| TY: 20-7-112 [16] | 70M | 79.6% | 0.91M | 53% | +30% |
| TY: 20-3-224 [16] | 218M | 36.4% | 3.34M | 23% | +0% |
| MCUNet [13] | 168M | 51.0% | 1.2M | 51.4% | +28.4% |
| MCUNetV2-M4 [12] | 172M | 49.9% | 0.47M | 64.6% | +41.6% |
| MCUNetV2-H7 [12] | 343M | 0% | 0.67M | 68.3% | +45.3% |
| ELASTIC (OURS) | 86M | 74.9% | 1.36M | 72.3% | +49.3% |
A key practical challenge in ELASTIC’s outer-loop cyclic module-wise optimization is maintaining optimization continuity when switching the active module (e.g., backbone neck/head backbone). Although the inner-loop evolutionary optimizer is standard, naive alternation effectively cold-starts each stage by discarding the module-specialized population, forcing repeated re-discovery of candidates compatible with the fixed context.
To make alternation workable, we introduce Population Passthrough, a lightweight memory mechanism that transfers elite architectural “knowledge” across module switches. For each module (backbone vs. neck/head), we maintain a passthrough buffer containing the top-performing candidates (ranked by validation mAP under the current resource constraints) from the most recent optimization of . When ELASTIC revisits , we initialize the population by mixing a fixed fraction of elites from (inheritance) with newly sampled/mutated candidates (diversity), balancing exploitation and exploration.
This mechanism is new in our setting because it targets alternating modular NAS: unlike standard elitism/aging schemes that assume a single, fixed search space and continuous run, Population Passthrough is applied at module boundaries to stabilize the outer-loop alternation and avoid repeated re-discovery.
III-D Resource Allocation and Search Budget
To ensure deployability on resource-constrained hardware, we explicitly partition the overall memory budget between architectural modules. Let and denote the memory allocations for the backbone and head, respectively, such that . This modular budget decomposition is critical for embedded systems, where hardware constraints—such as on-chip SRAM, flash storage, and runtime latency—are rigid and non-negotiable [17]. By enforcing per-module constraints during search, ELASTIC ensures that all discovered subnetworks are compatible with the deployment requirements of target TinyML platforms.
In addition, ELASTIC is designed to support search interruption, allowing the iterative process to be paused at any stage and a valid, high-quality subnetwork to be extracted. This property is particularly beneficial in real-world TinyML development workflows, where compute resources are limited and iterative refinement may need to occur incrementally [11].
IV Experiments
| Dataset | Search Space | Mean mAP | mAP | Std. Dev. | Variance |
| SVHN | Joint Search [8] | 44.61% | 0% | 1.39 | |
| Neck/Head Search [7, 10, 24] | 67.38% | +22.77% | 0.37 | 0.136 | |
| ELASTIC(3) | 70.42% | +25.81% | 0.47 | 0.221 | |
| ELASTIC(5) | 70.68% | +26.07% | 0.45 | 0.199 | |
| PascalVOC | Joint Search [8] | 4.99% | 0% | 0.04 | |
| Neck/Head Search [7, 10, 24] | 27.89% | +22.90% | 0.01 | ||
| ELASTIC(3) | 28.28% | +23.29% | 0.01 | ||
| ELASTIC(5) | 28.32% | +23.33% | 0.01 |
IV-A Experimental Setup
We construct five OFA-based supernets tailored for diverse TinyML-relevant datasets: (i) SVHN, (ii) VGGFace2, (iii) a five-class subset of PascalVOC, (iv) the full PascalVOC dataset, and (v) TrashNet. These datasets are selected to cover a range of TinyML application domains, from lightweight digit recognition to face identification and general-purpose object detection, while avoiding the prohibitive memory and compute demands of large-scale datasets such as COCO.
The SVHN, VGGFace2, and PascalVOC subset supernets adopt a one-stage SSD-style architecture, while the full PascalVOC and TrashNet supernets adopt a MobileNet–YOLO style design that combines a lightweight MobileNet backbone with YOLO-style detection heads. For the SSD-style supernets, the search space varies (i) layer widths, (ii) kernel sizes (, ), and (iii) depth, i.e., the number of layers allocated at each feature map. For both the MobileNet–YOLO supernets, the search space varies (i) expansion ratios, (ii) kernel sizes (, ), and (iii) depth, i.e., the number of blocks allocated at each feature map.
All supernets are trained with a batch size of 64, learning rate of 0.001, momentum of 0.9, and weight decay of . Training follows the progressive shrinking strategy [2], with 200 epochs per level transition. Experiments are conducted on a single NVIDIA RTX 4090 GPU.
We evaluate deployment across three MCU platforms. The MAX78000/02 are accelerator-equipped microcontrollers with a dedicated CNN engine and separate on-chip memory for weights and activations. MAX78000 provides 442 kB weight memory and 524 kB data memory for the earliest streaming layers, while subsequent layers must fit within a single 32 kB memory instance. MAX78002 provides 2.4 MB weight memory and 1.3 MB data memory for the earliest streaming layers, while subsequent layers must fit within a single 80 kB memory instance. In contrast, the STM32F746 is a general-purpose microcontroller without a CNN accelerator, with 320 kB SRAM and 1 MB flash, where weights are stored in flash and activations are stored in SRAM. In this way, MAX78000/MAX78002 represent accelerator-equipped TinyML targets, while STM32F746 provides a non-accelerator MCU baseline for comparison.
Evolutionary search is performed with a population size of 100, mutation probability of 0.2, mutation ratio of 0.5, and a parent selection ratio of 0.25. Each search is budgeted for 55 generations on SVHN and VGGFace2, and 30 generations on the PascalVOC subset, full PascalVOC, and TrashNet datasets, respectively. Convergence is defined as the point at which the best candidate architecture’s mAP is within 1% of the final value achieved at the end of the search.
IV-B Comparative Search Evaluation
We evaluate the effectiveness of ELASTIC by comparing it against three baseline NAS strategies: backbone-only search [4], head-only search [22], and progressive NAS [22]. All methods are evaluated under a fixed budget using a consistent once-for-all supernet design.
Results on SVHN and PascalVOC subset.
Table I summarizes our search results. On SVHN, ELASTIC achieves a final mAP of 80.09%, outperforming progressive NAS (79.62%) while using only 12.5 GPU hours—a 2 speedup compared to the 30.8 hours required by progressive search. On PascalVOC, the improvement is more substantial: ELASTIC reaches 30.83% mAP, outperforming progressive NAS (21.74%) by +9.09 percentage points with similar compute (14.65 vs. 14.78 GPU hours).
Effect of Decoupling.
To isolate the impact of coordinated multi-module optimization, we compare ELASTIC against decoupled strategies that search only the backbone or the head. On SVHN, these methods achieve significantly lower performance: backbone-only search yields 75.53% mAP and head-only search reaches just 75.35%, falling short of ELASTIC’s 80.09% by 4.75 percentage points.
Convergence Behavior.
Figure 2 shows that ELASTIC converges up to 2 faster than progressive NAS. On SVHN, ELASTIC converges in just 15 iterations, compared to over 30 iterations for progressive NAS. In contrast, backbone-only and head-only searches stagnate early, plateauing at 75.53% and 75.35% mAP, respectively. A similar trend holds on PascalVOC, where ELASTIC rapidly reaches 30.83% mAP, outperforming progressive NAS (21.74%) and single-module baselines (22.02% and 19.22%). This disparity is exacerbated by progressive NAS’s rigid search pipeline, which fixes early module choices and prevents adaptation to downstream module configurations.
Comparison with MCUNET and TinyissimoYOLO.
To assess ELASTIC’s suitability for microcontrollers, we extend our experiments using the MCUNET [13] search space with the neck module included for fair comparison. Benchmarked on the full PascalVOC () dataset with similar parameter budgets (Table II), our ELASTIC-derived detector achieves 72.3% mAP, surpassing MCUNET by 20.9%, TinyissimoYOLO’s [15] best model by 16.3%, and MCUNetV2-M4/H7 [12] by 7.7% and 4.0%, respectively, while using 50–75% fewer MACs. Unlike MCUNetV2 [12], which requires external patch-based inference, ELASTIC supports native end-to-end inference, making it easier to deploy across diverse microcontrollers.
| Model | Dataset | Device | Params | Energy (J) | Latency (ms) | Power (mW) | mAP (%) |
| ai87-fpndetector [9] | PascalVOC | MAX78002 | 2.18M | 62001 | 122.6 | 445.76 | 50.66 |
| ELASTIC (OURS) | PascalVOC | MAX78002 | 1.32M | 17581 | 51.1 | 285.02 | 57.65 |
| ai87-fpndetector [9] | PascalVOC | STM32746 | 0.82M | N/A | 494.8 | N/A | 9.12 |
| ELASTIC (OURS) | PascalVOC | STM32746 | 0.89M | N/A | 490.6 | N/A | 19.0 |
| ai87-fpndetector [9] | TrashNet | MAX78002 | 2.18M | 62001 | 122.6 | 445.76 | 83.1 |
| ELASTIC (OURS) | TrashNet | MAX78002 | 1.32M | 17581 | 51.1 | 285.02 | 93.3 |
| ai87-fpndetector [9] | TrashNet | STM32746 | 0.82M | N/A | 494.8 | N/A | 54.03 |
| ELASTIC (OURS) | TrashNet | STM32746 | 0.89M | N/A | 490.6 | N/A | 73.9 |
| ai85net-tinierssd-face [9] | VGGFace2 | MAX78000 | 0.28M | 1712 | 43.4 | 29.90 | 84.72 |
| ELASTIC (OURS) | VGGFace2 | MAX78000 | 0.22M | 1368 | 45.6 | 20.90 | 87.10 |
| ai85net-tinierssd-face [9] | VGGFace2 | STM32746 | 0.277M | N/A | 968.2 | N/A | 80.2 |
| ELASTIC (OURS) | VGGFace2 | STM32746 | 0.168M | N/A | 699.1 | N/A | 82.1 |
| ai85net-tinierssd [9] | SVHN | MAX78000 | 0.19M | 573 | 14.0 | 29.20 | 83.60 |
| ELASTIC (OURS) | SVHN | MAX78000 | 0.22M | 341 | 13.0 | 16.70 | 88.10 |
| ai85net-tinierssd [9] | SVHN | STM32746 | 0.229M | N/A | 449.3 | N/A | 83.5 |
| ELASTIC (OURS) | SVHN | STM32746 | 0.235M | N/A | 523.7 | N/A | 87.9 |
IV-C Impact of Population Passthrough on Search Stability
Iterative NAS can suffer sharp drops at module boundaries because switching the active module effectively resets the evolutionary population. Without population memory, each alternation incurs a repeated “catch-up” phase to re-discover compatible high-quality candidates, harming stability and slowing convergence. We mitigate this with Population Passthrough, which carries a fixed fraction of elite architectures into the next stage to preserve optimization momentum.
Figure 3 ablates the passthrough ratio on PascalVOC. With 0% passthrough, mAP collapses after switches and saturates below 22%. Ratios of 50–80% stabilize the search and reach mAP in roughly half the iterations. 60% performs best, balancing exploitation (stability) and exploration (progress).
IV-D Iterative Search Space Refinement
To evaluate whether ELASTIC improves the effective search space, we randomly sample 100 architectures from three matched spaces: (1) joint backbone–neck/head, (2) progressive head-only (fixed backbone), and (3) ELASTIC-refined head-only (after re-optimization) in Figure 4. On SVHN, the mean mAP increases from 44.61% (joint) to 67.38% (progressive) and to 72.25% with ELASTIC (+4.87% vs. progressive). On PascalVOC, the mean improves from 4.99% (joint) to 27.89% (progressive) and to 28.32% with ELASTIC.
We further track refinement over iterations by repeating the sampling at iterations 0, 3, 5 (Fig. 5, Table III), treating iteration 0 as the progressive baseline. On SVHN, the mean mAP rises from 67.38% (iter. 0) to 70.68% (iter. 5), with variance changing from 0.136 to 0.199. On PascalVOC, the mean increases from 27.89% to 28.32%, while variance decreases from to . Overall, iterative revisiting yields a stronger induced space than progressive partitioning and improves stability on PascalVOC.
IV-E Deployment on MCU Platforms
The MAX78000/MAX78002 MCUs integrate a CNN accelerator but impose tight constraints on kernel sizes, layer counts, and memory. Despite these limits, ELASTIC-derived models improve the efficiency–accuracy trade-off over TinierSSD/FPN baselines [9] while remaining within on-chip budgets.
We highlight these results in Table IV. On PascalVOC, energy decreases by 71.6%, latency by 2.40, power by 36.1%, and mAP increases by +6.99 on MAX78002. On TrashNet, ELASTIC follows the same pattern on MAX78002: energy decreases by 71.6%, latency by 2.40, power by 36.1%, and mAP increases by +10.2. ELASTIC also generalizes across tasks: on VGGFace2, it reduces energy and power by 20.1% and 30.1%, respectively, while improving mAP by +2.38; on SVHN, it reduces energy and power by 40.5% and 42.8%, respectively, with a +4.50 mAP gain and a slight latency improvement. We further evaluate cross-device portability on STM32F746. On PascalVOC, the scaled ELASTIC model reduces latency slightly from 494.8 ms to 490.6 ms while improving mAP from 9.12 to 19.0. On TrashNet, the scaled ELASTIC model reduces latency slightly from 494.8 ms to 490.6 ms while improving mAP from 54.03 to 73.9. On VGGFace2, ELASTIC reduces latency by 27.8% and improves mAP by +1.9. On SVHN, ELASTIC improves mAP by +4.4.
Since the two scaled STM32F746 models differ in their maximum deployable input resolution, we also compare them at both and in Table V. The scaled ai87-fpndetector baseline can be deployed only up to , while the scaled ELASTIC model remains deployable at ; at , the baseline encounters out-of-memory due to activation SRAM limits on STM32F746. Under matched resolution, ELASTIC still outperforms the baseline on both datasets, improving mAP from 9.12 to 19.0 on PascalVOC and from 54.03 to 73.9 on TrashNet, while maintaining nearly identical latency (490.6 ms vs. 494.8 ms). Moreover, ELASTIC supports the higher resolution and achieves 40.1 mAP on PascalVOC and 86.0 mAP on TrashNet, further highlighting that the searched architecture is not only more accurate but also more deployable under tight STM32F746 memory constraints.
Figure 6 compares ELASTIC architectures to scaled TinierSSD baselines [9] and shows ELASTIC attains higher accuracy at lower compute (66% FLOPs reduction), underscoring the value of hardware-aware NAS for MCU deployment.
| Model | Dataset | Device | Params | Resolution | Latency (ms) | mAP (%) |
| ai87-fpndetector [9] | PascalVOC | STM32746 | 0.82M | 128128 | N/A | OOM |
| ai87-fpndetector [9] | PascalVOC | STM32746 | 0.82M | 100100 | 494.8 | 9.12 |
| ELASTIC (OURS) | PascalVOC | STM32746 | 0.89M | 128128 | 825.55 | 40.1 |
| ELASTIC (OURS) | PascalVOC | STM32746 | 0.89M | 100100 | 490.6 | 19.0 |
| ai87-fpndetector [9] | TrashNet | STM32746 | 0.82M | 128128 | N/A | OOM |
| ai87-fpndetector [9] | TrashNet | STM32746 | 0.82M | 100100 | 494.8 | 54.03 |
| ELASTIC (OURS) | TrashNet | STM32746 | 0.89M | 128128 | 825.55 | 86.0 |
| ELASTIC (OURS) | TrashNet | STM32746 | 0.89M | 100100 | 490.6 | 73.9 |
V Conclusion
We introduced ELASTIC, a unified framework for constrained, once-for-all NAS tailored to object detection on resource-limited platforms. By alternating backbone and head optimization, ELASTIC enables cross-module co-adaptation while significantly reducing search cost. A key component of our approach, Population Passthrough, stabilizes module transitions and accelerates convergence, addressing a major limitation of existing progressive and decoupled NAS methods.
Through extensive experiments on multiple datasets and deployment on real microcontroller platforms, we demonstrated that ELASTIC discovers architectures that are more accurate and more efficient than manually designed and standard NAS baselines. These results highlight the effectiveness of iterative, hardware-aware NAS in advancing the accuracy–efficiency frontier for resource-constrained edge devices.